34. A set of n 5 pairs of X and Y scores has Σχ = 15, ΣΥ = 5, and ΣΧΥ = 10. For these data, the value of SP is- a. c.-10 d. 10 b. 5
Answer #1

we know that SP =\sum XY -(\sum X\sum Y)/n

= 10 - 75/5

=10-15

= -5

hence A is right answer